# Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance in vision-language tasks, but their application in the medical field remains underexplored, particularly for integrating structured time series data with unstructured clinical notes. In clinical practice, dynamic time series data such as lab test results capture critical temporal patterns, while clinical notes provide rich semantic context. Merging these modalities is challenging due to the inherent differences between continuous signals and discrete text. To bridge this gap, we introduce ProMedTS, a novel self-supervised multimodal framework that employs prompt-guided learning to unify these heterogeneous data types. Our approach leverages lightweight anomaly detection to generate anomaly captions that serve as prompts, guiding the encoding of raw time series data into informative embeddings. These embeddings are aligned with textual representations in a shared latent space, preserving fine-grained temporal nuances alongside semantic insights. Furthermore, our framework incorporates tailored self-supervised objectives to enhance both intra- and inter-modal alignment. We evaluate ProMedTS on disease diagnosis tasks using real-world datasets, and the results demonstrate that our method consistently outperforms state-of-the-art approaches.
Unlocking Multimodal Integration in EHRs: A Prompt Learning Framework for Language and Time Series Fusion
Shuai Niu1, Jing Ma1, Hongzhan Lin1, Liang Bai2, Zhihua Wang3, Wei Bi5, Yida Xu1, Guo Li4, and Xian Yang6thanks: * This is the corresponding author. Hong Kong Baptist University1, Shanxi University2, Shanghai Institute for Advanced Study of Zhejiang University3, Manchester Metropolitan University4, Tencent AI Lab5, The University of Manchester6 {cssniu,majing, cshzlin}@comp.hkbu.edu.hk, sxbailiang@126.com, zhihua.wang@zju.edu.cn, victoriabi@tencent.com, xuyida@hkbu.edu.hk, l.guo@mmu.ac.uk, xian.yang@manchester.ac.uk
@@ 1 Introduction
Recent advancements in natural language processing (NLP) have revolutionized healthcare by enabling deeper insights into electronic health records (EHRs). EHRs combine structured data, such as time series laboratory (lab) test results, with unstructured data, including clinical notes and medical images. While large language models (LLMs) excel at processing unstructured text Nori et al. (2023); Singhal et al. (2023) and vision transformers have driven progress in medical image analysis Wang et al. (2022); Chen et al. (2021), integrating continuous time series data with text remains a challenge. Unlike text, which is composed of discrete tokens, time series data contain continuous signals with temporal dependencies Jin et al. (2023) as illustrated in Figure 1(a).
@@
Current multimodal learning approaches, especially contrastive learning methods Radford et al. (2021); Li et al. (2023), have been effective in aligning vision and language. However, they are less suited to bridge the gap between time series and text. Time series data require fine-grained temporal representations in a high-dimensional space and are often irregularly sampled, exhibit diverse frequencies, and include missing values Harutyunyan et al. (2019a). In addition, the lack of large-scale paired datasets that link raw time series with textual descriptions further hampers LLMs from incorporating structured information into clinical decision-making Niu et al. (2024). Without an effective fusion mechanism, LLMs cannot fully exploit the rich temporal patterns in structured EHR data.
@@
To address these challenges, we propose ProMedTS, a self-supervised and prompt-guided framework designed to unify medical notes and time series lab test for naturally understood by LLMs. As shown in Figure 1(b), instead of feeding raw time series directly into LLMs, our framework uses anomaly descriptions, capturing key patterns in lab test results for helping multimodal fusion between text and time series data. These descriptions are generated using lightweight anomaly detection techniques Vinutha et al. (2018) and convert continuous signals into human-readable summaries. The process involves two steps. First, anomaly descriptions establish a direct connection between time series EHRs and medical notes. Second, time series prompt embeddings are generated and added as prefix tokens to the LLM input. This method integrates structured time series information into the language modeling process without altering the LLM architecture, unifying both modalities within a same encoding space and enhancing clinical decision-making.
@@
We optimize ProMedTS with three self-supervised learning objectives. A contrastive loss maps textual and time series modalities into a shared latent space. An anomaly-time series matching loss links lab test with their corresponding anomaly descriptions to reinforce consistency. Finally, an anomaly caption generation loss improves the fine-grained alignment between numeric time series lab test and time series prompt embeddings. Together, these objectives enable LLMs to process both structured and unstructured EHR data more effectively, addressing the gap between language and time series representations in healthcare applications.
@@ 2 Related Work
The increasing diversity of EHR data has led to significant advancements in multimodal learning for healthcare applications. MedCLIP Wang et al. (2022) employs semantic contrastive learning to align medical images with textual reports, while RAIM Qiao et al. (2019) and GLoRIA Huang et al. (2021) integrate numerical and image data with text using attention mechanisms. LDAM Niu et al. (2021a) further extends these approaches by leveraging cross-attention with disease labels to fuse features from lab tests and clinical notes. EHR-KnowGen Niu et al. (2024) transforms structured lab data into text and incorporates external knowledge for improved modality fusion. Despite these advancements, achieving a unified latent embedding that effectively captures interactions across diverse modalities remains a key challenge in multimodal EHR processing.
Beyond multimodal learning, recent research has explored generative approaches to healthcare modeling. Conventional methods have primarily relied on discriminative models for disease risk assessment and diagnosis Choi et al. (2016); Niu et al. (2021b); Qiao et al. (2019). However, generative models are increasingly being adopted, as demonstrated by Clinical CoT Kwon et al. (2024), applying LLMs for disease diagnosis generation. Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) and Chain-of-Thought (CoT) prompting Wei et al. (2022) have further enhanced medical reasoning capabilities in models such as GatorTron Yang et al. (2022), MedPalm Singhal et al. (2023), and GPT4-Med Nori et al. (2023). While these models excel in medical question-answering, they remain limited in real-world direct disease diagnosis and multimodal EHR processing. EHR-KnowGen Niu et al. (2024) reframes disease diagnosis as a text-to-text generation problem but overlooks the crucial temporal details embedded in time series lab tests, underscoring the need for more effective and dedicated multimodal fusion strategies.
@@ 3.1 Problem Definition
We introduce ProMedTS, aiming to reduce discrepancies between language and time series data in EHRs. Specifically, it leverages anomaly captions and generates time series prompt embeddings to unify both modalities in a shared latent space. The inputs to ProMedTS, denoted by $\{\bm{M},\bm{X}\}$, include medical notes $\bm{M}\in\mathbb{R}^{B\times N_{m}}$ (where $B$ is the batch size and $N_{m}$ is the number of tokens) and numeric lab test data $\bm{X}\in\mathbb{R}^{B\times L\times N_{x}}$ (where $L$ is the sequence length and $N_{x}$ is the number of lab test variants). Additionally, a lightweight anomaly detection Vinutha et al. (2018) is employed to generate textual descriptions of anomalies $\bm{C}\in\mathbb{R}^{B\times N_{c}}$ (details in Appendix A.2). ProMedTS also uses learnable time series query embeddings $\bm{P}\in\mathbb{R}^{B\times N_{p}\times D}$, which are transformed into time series prompt embeddings $\bm{\mathcal{T}}\in\mathbb{R}^{B\times N_{p}\times D}$, where $N_{p}$ is the query length and $D$ is the hidden dimension.
@@ 3.2 Model Overview
Figure 2 illustrates the overview of ProMedTS, which comprises three main modules. Three modules share the same Clinical-BERTAlsentzer et al. (2019) structured model and are extended to support cross-attention, self-attention, and prompt generation. The Time Series Prompt Embedding (TSPE) module applies a cross-attention mechanism to convert raw lab test data into prompt embeddings, preserving key temporal features. The Multimodal Textual Information Fusion (MTIF) module encodes and merges medical notes with anomaly captions in a unified latent space, facilitating the extraction of complementary semantic information. Finally, the Self-supervised Learning (SSL) module employs tailored loss functions to bridge the modality gap and maintain fine-grained temporal details in the learned representations. These modules work in tandem to achieve robust alignment and fusion of heterogeneous EHRs, and the following sections provide in-depth explanations of each component and their applications.
@@ 3.3 Time Series Prompt Embedding
The objective of the TSPE module is to extract and encapsulate the inherent information from time series lab test data into a time series prompt embedding. Let $\{\bm{X},\bm{P}\}$ represent the module inputs. The numeric lab test data $\bm{X}$ is first processed by a time series encoder (TSE) using PatchTST Nie et al. (2022). In parallel, the learnable query embeddings $\bm{P}$,initialized using vectors extracted from the Clinical-BERT word embedding layer, serve as query tokens in the cross-attention mechanism, guiding the selection of relevant temporal features by attending to time series lab test results encoded by TSE. To generate the final prompt embedding $\bm{\mathcal{T}}$, we extend the multi-head self-attention encoder of Clinical-BERT to support a multi-head cross-attention mechanism, following a strategy similar to that adopted in Li et al. (2023). We designate $\bm{X}$ as both key and value while $\bm{P}$ serves as the query:
@@ 3.4 Multimodal Textual Information Fusion
where $\bm{E}_{m}\in\mathbb{R}^{B\times N_{m}\times D}$ and $\bm{E}_{c}\in\mathbb{R}^{B\times N_{c}\times D}$. The repeated inputs indicate that the key, query, and value matrices are identical for the self-attention mechanism. This structure enables the model to encode each type of textual information independently while capturing the inherent characteristics and context of each input. The combined textual representation is then derived from these encoded inputs:
@@ 3.5 Self-Supervised Learning
This module addresses the modality gap between textual and time series EHR data using three specialized loss functions. By simultaneously aligning cross-modal representations and preserving fine-grained temporal details, the model learns to capture both semantic and temporal nuances.
@@ 3.5.1 Cross-Modal Contrastive Alignment
To promote cross-modal alignment, we design a contrastive loss that brings language and time series embeddings closer when they originate from the same patient and pushes them apart otherwise. We first compute similarity matrices by multiplying the fused text representation $\bm{E}_{f}$ with the time series prompt embeddings $\bm{\mathcal{T}}$:
@@
where the $\max$ operator performs max-pooling across $N_{p}$ dimensions, yielding $\bm{g}_{(\bm{E}_{f},\bm{X})}$ and $\bm{g}_{(\bm{X},\bm{E}_{f})}\in\mathbb{R}^{B\times B}$. Note that $\bm{g}_{(\bm{E}_{f},\bm{X})}$ measures text-to-time series similarity (by fixing $\bm{E}_{f}$ and iterating over $\bm{\mathcal{T}}$), while $\bm{g}_{(\bm{X},\bm{E}_{f})}$ captures time-series-to-text similarity (by fixing $\bm{\mathcal{T}}$ and iterating over $\bm{E}_{f}$). We then apply the SoftMax function to generate two distinct sets of logits:
@@ 3.5.2 Intra-Modal Matching
To further capture intra-modality consistency, we align lab tests with corresponding anomaly descriptions. This alignment is modeled as a binary classification task, distinguishing matched from unmatched pairs of lab tests and anomaly captions. Following Li et al. (2021), we employ a negative mining strategy to generate labels $\bm{y}_{m}$ by selecting the most similar pairs in a training batch as negative samples, where the top 1-ranked pair is labeled as 1 and the others as 0, based on the similarity computed in Equation 4. We employ Clinical-BERT’s cross-attention, where the concatenation of $\bm{C}$ and $\bm{P}$ serves as the query, and the encoded time series $\bm{X}$ is used as both key and value. A Multilayer Perceptron (MLP) classifier with softmax activation, denoted $f_{match}$, predicts the probability $\hat{\bm{y}}_{m}$:
@@ 3.5.3 Anomaly Description Reconstruction
This objective encourages the model to generate accurate textual descriptions, thereby reinforcing alignment between time series prompts and language tokens.
@@ 3.6 LLM-based Disease Diagnosis with ProMedTS
To illustrate the practical effectiveness of ProMedTS in unifying textual and time series data, we employ a pre-trained, frozen LLM model for disease diagnosis. As depicted in Figure 3, ProMedTS first converts numeric lab test results into time series prompt embeddings, which are then aligned via a fully connected layer to match the LLM’s input dimensions. These embeddings serve as prefix soft prompts, concatenated with the medical notes so that the model can ingest structured signals from time series alongside unstructured clinical text. By bridging language and time series modalities, the LLM can process both inputs concurrently, leveraging complementary information for enhanced diagnostic accuracy.
@@
ProMedTS 267.5M 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
ProMedTS* 1B 60.62(0.22) 67.83(0.18) 64.02(0.11) 59.43(0.37) 63.65(0.54) 60.78(0.13)
@@
ProMedTS 267.5M 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
ProMedTS* 1B 71.12(0.31) 69.33(0.42) 70.21(0.05) 70.97(0.42) 65.51(0.64) 67.56(0.09)

### v2

# ProMedTS: A Self-Supervised, Prompt-Guided Multimodal Approach for Integrating Medical Text and Time Series

@@ Abstract
Large language models (LLMs) have shown remarkable performance in vision-language tasks, but their application in the medical field remains underexplored, particularly for integrating structured time series data with unstructured clinical notes. In clinical practice, dynamic time series data, such as lab test results, capture critical temporal patterns, while clinical notes provide rich semantic context. Merging these modalities is challenging due to the inherent differences between continuous signals and discrete text. To bridge this gap, we introduce ProMedTS, a novel self-supervised multimodal framework that employs prompt-guided learning to unify these heterogeneous data types. Our approach leverages lightweight anomaly detection to generate anomaly captions that serve as prompts, guiding the encoding of raw time series data into informative prompt embeddings. These prompt embeddings are aligned with textual representations in a shared latent space, preserving fine-grained temporal nuances alongside semantic insights. Furthermore, our framework incorporates tailored self-supervised objectives to enhance both intra- and inter-modal alignment. We evaluate ProMedTS on disease diagnosis tasks using real-world datasets, and the results demonstrate that our method consistently outperforms state-of-the-art approaches.
@@ 1 Introduction
Recent advancements in natural language processing (NLP) have revolutionized healthcare by enabling deeper insights into electronic health records (EHRs). EHRs combine structured data, such as time series laboratory (lab) test results, with unstructured data, including clinical notes and medical images. While large language models (LLMs) excel at processing unstructured text Nori et al. (2023); Singhal et al. (2023) and vision transformers have driven progress in medical image analysis Wang et al. (2022); Chen et al. (2021), integrating time series data with text remains a challenge. Unlike text, which is composed of discrete tokens, time series data contains continuous signals with temporal dependencies as illustrated in Figure 1(a).
@@
Current multimodal learning approaches, especially contrastive learning methods Radford et al. (2021); Li et al. (2023), have been effective in aligning vision data and text data. However, they are less suited to bridge the gap between time series data and text data. Time series data require fine-grained temporal representations in a high-dimensional space and are often irregularly sampled, exhibit diverse frequencies, and include missing values Harutyunyan et al. (2019a). In addition, the lack of large-scale paired datasets that link raw time series with textual descriptions further hampers LLMs from incorporating structured information into clinical decision-making Niu et al. (2024a). Without an effective fusion mechanism, LLMs cannot fully exploit the rich temporal patterns in structured EHR data.
@@
To address these challenges, we propose ProMedTS, a self-supervised and prompt-guided framework designed to unify medical notes and time series lab tests for natural comprehension by LLMs. As illustrated in Figure 1(b), rather than only feeding raw time series data directly into LLMs, our framework introduces anomaly descriptions as a modality bridge to facilitate alignment between text and time series data and to aid in generating time series prompt embeddings. These descriptions are produced using anomaly detection technology Vinutha et al. (2018), which converts continuous signals into human-readable summaries that encompass coarse-grained time series patterns. The process consists of two steps. First, anomaly descriptions establish a direct connection between time series lab tests and medical notes. Second, the generated time series prompt embeddings contain both coarse-grained anomaly information and fine-grained time series variation patterns. These embeddings are then appended as prefix tokens to the input of the LLM. This approach integrates structured time series information into the language modeling process without modifying the LLM architecture, thereby unifying both modalities within the same encoding space and enhancing clinical decision-making.
@@
We optimize ProMedTS with three self-supervised learning objectives for learning the time series prompt embeddings. A contrastive loss maps textual and time series modalities into a shared latent space to learn the coarse-grained lab test information. An anomaly-time series matching loss links numeric lab tests with their corresponding anomaly descriptions to reinforce consistency and learn the fine-grained lab test information. Finally, an anomaly caption generation loss enhances the representation of multi-granularity time series information within the time series prompt embeddings. Together, these objectives enable LLMs to process both structured and unstructured EHR data more effectively, addressing the gap between language and time series representations in healthcare applications.
@@ 2 Related Work
2.1 Multimodal Learning in Healthcare
The increasing diversity of EHR data has led to significant advancements in multimodal learning for healthcare applications. MedCLIP Wang et al. (2022) employs semantic contrastive learning to align medical images with textual reports, while RAIM Qiao et al. (2019) and GLoRIA Huang et al. (2021) integrate numerical or image data with text using attention mechanisms. LDAM Niu et al. (2021) further extends these approaches by leveraging cross-attention with disease labels to fuse features from lab tests and clinical notes. EHR-KnowGen Niu et al. (2024a) transforms structured lab data into text and incorporates external knowledge for improved modality fusion. Despite these advancements, achieving a unified latent embedding that effectively captures interactions across diverse modalities remains a key challenge in multimodal EHR processing.
2.2 Healthcare with LLMs
Beyond multimodal learning, recent research has explored generative approaches to healthcare modeling. Conventional methods have primarily relied on discriminative models for disease risk assessment and diagnosis Choi et al. (2016); Niu et al. (2024b); Qiao et al. (2019). However, generative models are increasingly being adopted, as demonstrated by Clinical CoT Kwon et al. (2024), applying LLMs for disease diagnosis generation. Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) and Chain-of-Thought (CoT) prompting Wei et al. (2022) have further enhanced medical reasoning capabilities in models such as GatorTron Yang et al. (2022), MedPalm Singhal et al. (2023), and GPT4-Med Nori et al. (2023). While these models excel in medical question-answering, they remain limited in real-world direct disease diagnosis and multimodal EHR processing. EHR-KnowGen Niu et al. (2024a) reframes disease diagnosis as a text-to-text generation problem but overlooks the crucial temporal details embedded in time series lab tests, underscoring the need for more effective and dedicated multimodal fusion strategies.
@@ 3.1 Problem Definition
We introduce ProMedTS, aiming to reduce discrepancies between language and time series EHRs. Specifically, it leverages anomaly captions and generates time series prompt embeddings to unify both modalities in a shared latent space. The inputs to ProMedTS, denoted by $\{\bm{M},\bm{X}\}$, include medical notes $\bm{M}\in\mathbb{R}^{B\times N_{m}}$ (where $B$ is the batch size and $N_{m}$ is the number of tokens) and numeric lab test data $\bm{X}\in\mathbb{R}^{B\times L\times N_{x}}$ (where $L$ is the sequence length and $N_{x}$ is the number of lab test variants). Additionally, a lightweight anomaly detection Vinutha et al. (2018) is employed to generate textual descriptions of anomalies $\bm{C}\in\mathbb{R}^{B\times N_{c}}$ (details in Appendix A.2). ProMedTS also uses learnable time series query embeddings $\bm{P}\in\mathbb{R}^{B\times N_{p}\times D}$, which are transformed into time series prompt embeddings $\bm{\mathcal{T}}\in\mathbb{R}^{B\times N_{p}\times D}$, where $N_{p}$ is the query length and $D$ is the hidden dimension.
@@ 3.2 Model Overview
Figure 2 illustrates the overview of ProMedTS, which comprises three main modules. Three modules share the same Clinical-BERTAlsentzer et al. (2019) structured model for clinical tokens embedding and are extended to support cross-attention, self-attention, and prompt generation. The Time Series Prompt Embedding (TSPE) module applies a cross-attention mechanism to convert raw lab test data into prompt embeddings, preserving key temporal features. The Multimodal Textual Information Fusion (MTIF) module encodes and merges medical notes with anomaly captions in a unified latent space, facilitating the extraction of complementary semantic information. Finally, the Self-supervised Learning (SSL) module employs tailored loss functions to bridge the modality gap and maintain multi-granularity temporal details in the learned representations. These modules work in tandem to achieve robust alignment and fusion of heterogeneous EHRs, and the following sections provide in-depth explanations of each component and its applications.
@@ 3.3 Time Series Prompt Embedding
The objective of the TSPE module is to extract and encapsulate the inherent fine-grained temporal information from time series lab tests into time series prompt embeddings. Let $\{\bm{X},\bm{P}\}$ represent the module inputs. The numeric lab test data $\bm{X}$ is first processed by a time series encoder (TSE) using PatchTST Nie et al. (2022). In parallel, the learnable query embeddings $\bm{P}$, initialized using vectors slicing from the Clinical-BERT word embedding layer, serve as query tokens in the cross-attention mechanism, guiding the selection of relevant temporal features by attending to time series lab tests encoded by TSE. To generate the final prompt embedding $\bm{\mathcal{T}}$, we extend the multi-head self-attention encoder of Clinical-BERT to support a multi-head cross-attention mechanism, following a strategy similar to that adopted in Li et al. (2023). We designate $\bm{X}$ as both key and value while $\bm{P}$ serves as the query:
@@ 3.4 Multimodal Textual Information Fusion
where $\bm{E}_{m}\in\mathbb{R}^{B\times N_{m}\times D}$ and $\bm{E}_{c}\in\mathbb{R}^{B\times N_{c}\times D}$. The repeated inputs indicate that the key, query, and value matrices are identical for the self-attention mechanism. This structure enables the model to encode each type of textual information independently while capturing the inherent characteristics and context of each input. The combined textual representation is then derived from $\bm{E}_{m}$ and $\bm{E}_{c}$:
@@ 3.5 Self-Supervised Learning
This module addresses the modality gap between textual and time series EHR data using three specialized loss functions. By simultaneously aligning cross-modal representations and preserving coarse- and fine-grained temporal details, the model learns to capture both semantic and temporal nuances.
@@ 3.5.1 Cross-Modal Contrastive Alignment
To promote cross-modal alignment and preserve coarse-grained temporal information into time series prompt embeddings, we design a contrastive loss that brings multimodal textual embedding and time series embeddings closer when they originate from the same patient and pushes them apart otherwise. We first compute similarity matrices by multiplying the fused text representation $\bm{E}_{f}$ with the time series prompt embeddings $\bm{\mathcal{T}}$:
@@
where the $\max$ operator performs max-pooling across $N_{p}$ dimensions, yielding $\bm{g}_{(\bm{E}_{f},\bm{X})}$ and $\bm{g}_{(\bm{X},\bm{E}_{f})}\in\mathbb{R}^{B\times B}$. Note that $\bm{g}_{(\bm{E}_{f},\bm{X})}$ measures text-to-time series similarity (by fixing $\bm{E}_{f}$ and iterating over $\bm{\mathcal{T}}$), while $\bm{g}_{(\bm{X},\bm{E}_{f})}$ captures time-series-to-text similarity (by fixing $\bm{\mathcal{T}}$ and iterating over $\bm{E}_{f}$). This process is the same as that used in vision-language contrastive learning Radford et al. (2021); Li et al. (2023). We then apply the SoftMax function to generate two distinct sets of logits:
@@ 3.5.2 Intra-Modal Matching
To further capture intra-modality consistency and enhance alignment with fine-grained temporal information, we align lab tests with their corresponding anomaly descriptions and time series prompt embeddings. This alignment is modeled as a binary classification task, distinguishing matched from unmatched pairs of lab tests and anomaly captions. Following Li et al. (2021), we employ a negative mining strategy to generate labels $\bm{y}_{m}$ by selecting the most similar pairs in a training batch as negative samples, where the top 1-ranked pair is labeled as 1 and the others as 0, based on the similarity computed in Equation 4. We employ Clinical-BERT’s cross-attention, where the concatenation of $\bm{C}$ and $\bm{P}$ serves as the query (aiming to align the time series prompt embedding more closely with the textual encoding space during training), and the encoded time series $\bm{X}$ is used as both key and value. A Multilayer Perceptron (MLP) classifier with softmax activation, denoted $f_{match}$, predicts the probability $\hat{\bm{y}}_{m}$:
@@ 3.5.3 Anomaly Description Reconstruction
This objective is a standard language model generation loss, computed as cross-entropy between the predicted token distribution and the ground truth tokens, encouraging the model to generate accurate textual descriptions, thereby reinforcing alignment between time series prompts and language tokens.
@@ 3.6 LLM-based Disease Diagnosis with ProMedTS
To illustrate the practical effectiveness of ProMedTS in unifying textual and time series data, we employ a pre-trained, frozen LLM model for disease diagnosis. As depicted in Figure 3, during ProMedTS’s fine-tuning, it first transforms numeric lab test results into time series prompt embeddings, which are subsequently aligned with the LLM’s input dimensions through a tunable fully connected layer. These embeddings then serve as prefix soft prompts, concatenated with the medical notes so that the model can ingest structured signals from time series alongside unstructured clinical text. By bridging language and time series modalities, the LLM can process both inputs concurrently, leveraging complementary information for enhanced diagnostic accuracy.
@@
ProMedTS
w/ OPT 316M 54.57(0.14) 54.92(0.21) 54.74(0.09) 55.92(0.08) 52.15(0.13) 53.68(0.07)
w/ QWEN-2.5 685M 55.37(0.15) 58.97(0.18) 57.82(0.13) 56.45(0.14) 56.72(0.16) 55.97(0.13)
w/ Flan-T5-small 268M 61.32(0.54) 66.65(0.51) 63.67(0.08) 60.35(0.61) 61.62(0.71) 60.42(0.18)
w/ Flan-T5-large 1B 60.62(0.22) 67.83(0.18) 64.02(0.11) 59.43(0.37) 63.65(0.54) 60.78(0.13)
@@
ProMedTS
w/ OPT 316M 54.99(0.15) 55.20(0.19) 55.10(0.11) 56.81(0.10) 52.36(0.17) 53.23(0.10)
w/ QWEN-2.5 685M 48.65(0.25) 72.32(0.17) 58.14(0.16) 57.35(0.20) 66.23(0.13) 57.51(0.12)
w/ Flan-T5-small 268M 71.63(0.46) 67.81(0.85) 69.69(0.18) 70.12(0.47) 63.58(0.79) 66.21(0.17)
w/ Flan-T5-large 1B 71.12(0.31) 69.33(0.42) 70.21(0.05) 70.97(0.42) 65.51(0.64) 67.56(0.09)